The first question) Calculate the percentage atom economy of using copper oxide and sulfuric acid to make copper sulfate:

CuO(s) + H2SO4(aq) --> CuSO4(aq) + H2O(l)

Give your answer to 3 significant figures


second question) Calculate the atom economy to make copper nitrate:

CuCO3(s) + 2HNO3(aq) --> Cu(NO3)2(aq) + CO2(g) + H2O(l)

Give your answer to 3 significant figures




1
Expert's answer
2021-10-12T02:07:29-0400

"% atom economy= mass of the desired product" Question 1

 % atom economy= mass of the desired reactant/Total mass of all the reactant  

                                  

Mass of desired reactant (copper oxide) =76


Total mass of all the reactant (copper oxide and sulphuric acid) =76+98=174

 

% atom economy (copper oxide) =   76/174* 100= 43.7%

                                                   

% atom economy (sulfuric acid) = mass of desired reactant/Total mass of the reactant

Mass of desired reactant(sulphuric acid) =98

Total mass of all the reactant (copper oxide and sulphuric acid) =76+98=174

                                                       


% atom economy (sulfuric acid) =  98/174*100 = 56%

                                                        Question 2

% atom economy (copper nitrate) = mass of the desired product/ Total mass of the all reactants

                                                      

Mass of Cu(NO3)2= 187.6

Mass of CuCO3 = 123.6


Mass of HNO3 = 2(63.0) = 126


% atom economy (copper nitrate) =  187.6/(123.6+126­) *100 = 74.9%